OpenStudy (anonymous):

How would you solve this without a calculator? 2((squarerootof3)/2) - 6((squarerootof3)/3)

OpenStudy (anonymous):

for the first half of the equation, the 2's cancel out so you are left with sqrt(3). for the second half of the equation, the 6 and 3 cancel to make 2sqrt(3) so you are left with \[\sqrt{3}-2\sqrt{3}=-1\sqrt{3}\]

OpenStudy (anonymous):

since the have the same value under the square root, you can subtract them